An ice cream store made a profit of $35700 in 1990 and a profit of $85360 in 2008. Write an equation that can be used to predict the profit, y, in terms of the year, x. Let x=0 represent the year 1990.
• Predict the profit for the year 2011.
• What does the y-intercept represent in the context of this problem?

Answer by stanbon(75887)

An ice cream store made a profit of $35700 in 1990 and a profit of $85360 in 2008. Write an equation that can be used to predict the profit, y, in terms of the year, x. Let x=0 represent the year 1990.
----
You have 2 points relating profit to year:: (8,85360) and (0,35700)
----
Equation:
P(x) = (85360/35700)t + 35700
-------------
• Predict the profit for the year 2011.
P(11) = (85360/35700)(11)+35700 = $35726.30
-------------------------------
• What does the y-intercept represent in the context of this problem?
P(0) = 35700 is the profit in year 1990
---------------
Cheers,
Stan H.